### v2

# pnasresearcharticle

Abstract
The epidemiological dynamics of Mycoplasma pneumoniae are characterized by complex and poorly understood multiannual cycles, posing challenges for forecasting. Using Bayesian methods to fit a seasonally forced transmission model to long-term surveillance data from Denmark (1958–1995, 2010–2025), we investigate the mechanisms driving recurrent outbreaks of M. pneumoniae. The period of the multiannual cycles (predominantly approx. 5 years in Denmark) are explained as a consequence of the interaction of two time-scales in the system, one intrinsic and one extrinsic (seasonal). While it provides an excellent fit to shorter time series (a few decades), we find that the deterministic model eventually settles into an annual cycle, failing to reproduce the observed 4-5-year periodicity long-term. Upon further analysis, the system is found to exhibit transient chaos and thus high sensitivity to stochasticity. We show that environmental (but not purely demographic) stochasticity can sustain the multi-year cycles via stochastic resonance. The disruptive effects of COVID-19 non-pharmaceutical interventions (NPIs) on M. pneumoniae circulation constitute a natural experiment on the effects of large perturbations. Consequently, the effects of NPIs are included in the model and medium-term predictions are explored. Our findings highlight the intrinsic sensitivity of M. pneumoniae dynamics to perturbations and interventions, underscoring the limitations of deterministic epidemic models for long-term prediction. More generally, our results emphasize the potential role of stochasticity as a driver of complex cycles across endemic and recurring pathogens.
keywords:
Mycoplasma pneumoniae |||| Epidemiological dynamics |||| Dynamical systems |||| Chaos |||| Stochasticity
[…]
This manuscript was compiled on April 16, 2025
[…]
The bacterial pathogen Mycoplasma pneumoniae is simultaneously common and unusual. It lacks a cell wall and is among the smallest free-living bacteria (1). The infection it causes has been given names such as (primary) atypical pneumonia (2), cold pneumonia and walking pneumonia (3) due to its tendency to cause less severe disease compared with typical bacterial pneumonia. It is a very common respiratory pathogen, especially in children and adolescents, and accounts for 15-20% of all cases of pneumonia (4). Disease courses are typically mild, but severe outcomes do occur. About 5-10% of upper airway infections with M. pneumoniae progress to pneumonia, and there are likely many asymptomatic cases (5), suggesting a high degree of under-ascertainment. Extrapulmonary symptoms are increasingly recognized (6, 7) and include complications such as encephalitis and RIME/Stevens-Johnson Syndrome (8, 9), with M. pneumoniae being the most common infectious cause of the latter in children (10, 11). As many as 25% of M. pneumoniae cases develop some extrapulmonary symptoms (12), most commonly dermatological or gastrointestinal; sometimes cardiovascular (13) and neurological (14) complications are seen, the latter in as many as 1-10% of those M. pneumoniae cases which require hospitalization (15, 12, 16).
From an epidemiological dynamics perspective, a remarkable feature of M. pneumoniae are the observed complex outbreak cycles, most often with a 3-7 year period (17, 18, 19, 20). In Denmark (2), as well as England and Wales (21), 4-5 year cycles have generally been observed. This is in contrast to some classical, strongly immunizing childhood diseases, such as measles, which tend to exhibit simpler annual or biennial dynamics (22), although there are exceptions to that rule, e.g. rubella (23, 24). Despite being such a common pathogen, the mechanisms behind the complex epidemiological dynamics of M. pneumoniae remain poorly understood, complicating prediction and forecasting. In general, predictability is important for timely resource allocation and deployment of relevant screening and diagnostic testing schemes. When the year-to-year incidence exhibits wide variability, planning for outbreaks becomes simultaneously more important, and more difficult. Discerning and understanding patterns in this variability is thus paramount for effective planning, not least for M. pneumoniae, where severe outcomes are typically preventable if diagnosed and treated in a timely manner.
Since early 2020, the multiannual cycles of M. pneumoniae have been largely disrupted (25, 26). This is an example of the polymicrobial effects of the COVID-19-related non-pharmaceutical interventions (NPIs), observed as a disruption of circulation across many pathogens (27) but perhaps evidenced most strikingly by the apparent extinction of the Yamagata strain of influenza B virus (28, 29).
Over three years, M. pneumoniae circulation was drastically curtailed in many locations and Denmark is no exception. A similar multi-year delay in M. pneumoniae resurgence has recently been reported for Europe and Asia (30) as well as the United States (25, 31). Such large-scale perturbations effectively provide a “natural experiment”, demonstrating how the dynamics of M. pneumoniae responds to a prolonged disruption and the associated build-up of susceptible individuals, as well as aiding model parameter estimation.
Here, we use a mathematical model that incorporates seasonality, vital dynamics (births/deaths), immunity loss and the effects of NPIs to study recent (2010–2025) and historical (1958–1995) M. pneumoniae incidence data from Denmark. The simple deterministic model is found to capture the observed dynamics well over a few decades but eventually reverts to annual outbreaks, failing to reproduce the prominent multiannual cycles long-term. We show that the presence of environmental noise can indefinitely sustain the multi-year outbreak patterns. In doing so, we highlight the important role that fluctuations in contact patterns, environmental noise, and demographic processes plays when forecasting pathogens with complex dynamics such as M. pneumoniae.
Lessons from contemporary and historical data
We utilize two different data sets as correlates of M. pneumoniae incidence, both from Denmark (Statens Serum Institut, SSI). One has M. pneumoniae-positive PCR tests (and the raw number of tests performed) with a weekly reporting frequency and covers the period 2010-2025. This data set forms the primary basis for parameter estimation, since the test data is highly M. pneumoniae-specific and temporally well-resolved. Secondarily, we use a data set covering the period 1958-1995. Beginning in 1946, patient blood samples collected at hospitals and general practitioners for the diagnosis of Primary Atypical Pneumonia (later identified as M. pneumoniae) were sent to SSI to undergo the cold agglutinin (CA) reaction. This reaction, unfortunately, has only low to moderate sensitivity ($\sim 20-80\%$ (32, 2)) and moderate specificity ($\sim 82-92\%$ (32, 2)) for M. pneumoniae. However, beginning in 1958, CA-positive samples were also subjected to M. pneumoniae Complement Fixation (MpCF), which was found to have higher specificity (33, 34). We thus also treat those tests from the 1958-1995 data set which are simultaneously CA- and MpCF-positive as an incidence proxy on which we base (secondary) parameter estimates.
The duration of immunity following M. pneumoniae infection is known not to be life-long, but the precise duration is unknown, with estimates ranging from a few years to more than a decade (35, 2, 25). For this reason, we opted for an SIRS (Susceptible-Infected-Recovered-Susceptible) compartmental model which allows previously infected individuals to become susceptible, and thus potentially reinfected. The model includes seasonal modulation of the transmission rate and equal birth/death rates (which are assumed constant within each data fit, for simplicity). The duration of immunity was treated as a parameter to be estimated during fitting to data.
The influence of NPIs are implemented as a slowly-varying multiplicative effect on the transmission rate from March 2020 onward, representing a reduction in overall contact rates, the magnitude of which is an estimated parameter (see Supplementary Fig. S5).
We estimate a basic reproductive number of $R_{0}=1.8$ (95%CI [1.7, 2.0], in good agreement with ref. (36)), a mean duration of immunity of 11 years and a relative amplitude of seasonality in transmission of $0.19$. All estimated parameters and their uncertainties can be found in Table S1.
With a sinusoidal seasonality function [^1] We opted for a sinusoidal seasonality function for simplicity, but have also directly estimated the shape of the seasonality function from the data. The resulting seasonality function can be found in Supplementary Fig. S4., the model fits the 2010-2025 data well (with the partial exception of the 2017 season in which the apparent incidence was somewhat higher than modeled), see Fig. 1A. The historical 1958-1995 data, which is overall of much lower fidelity, is captured more loosely by the model, see Fig. 1B. Interestingly, here we had to allow for a slow change in the mean transmission rate and ascertainment (allowed to change only slightly every 4 years), otherwise the model would tend to settle into an annual pattern.
This leads us to a more general observation: the SIRS model with M. pneumoniae parameters is able to capture the patterns reasonably well over a few decades, but in the absence of perturbations, the modeled dynamics always decays into an annual pattern eventually. This is of course at odds with the ubiquity of complex multiannual cycles in M. pneumoniae (2, 17, 18, 19, 20, 21) and other SIRS-like epidemiological systems such as parvovirus B19 (37, 38, 39, 40, 41). In the next section we discuss how the time-scales of the M. pneumoniae system may give rise to an approximately 5-year cycle, at least transiently. We then go on to investigate the role that stochasticity plays in sustaining such cycles over decades and in re-igniting them even after periods of approximately annual dynamics. In the final part of the paper, we study projections of future M. pneumoniae incidence with and without stochasticity.
Figure 1: Model fits to Mycoplasma pneumoniae incidence data. A seasonal SIRS model is fitted to recent (2010-2025) and historical (1958-1995) M. pneumoniae testing data from Denmark using Bayesian inference. A) The seasonal SIRS model explains observed M. pneumoniae dynamics 2010-2025 well. The fitted data are PCR positivity rates. B) The same type of model captures key features of observed M. pneumoniae dynamics 1958-1995. The fitted data are variance-stabilized numbers of samples which are simultaneously CA and MpCF-positive. (Data source: Statens Serum Institut and (2)).
Two time-scales conspire to create complex cycles
The seasonally forced SIRS system with vital dynamics contains two comparable oscillatory time-scales: the one-year seasonal period and the time-scale set by the intrinsic epidemiological dynamics. To ascertain the latter time-scale, the corresponding autonomous system may be studied (that is, the SIRS system with vital dynamics but no seasonal forcing). This system will eventually reach an endemic equilibrium, but will oscillate as it approaches this equilibrium. A linearization of the autonomous system around its endemic equilibrium yields an intrinsic oscillation frequency as well as an exponential decay rate (42), given by a combination of the transmission rate $\beta$, immune waning rate $\delta$, recovery rate $\gamma$ and birth/death rate $\mu$. Using the parameter values estimated from the Bayesian fit to Danish data (2010-2025), the frequency is found to be $f_{1}=1/(43.46\text{wk})$ (95%CI: [1/44.0, 1/43.0]). This natural oscillatory time-scale offers a clue to how the 4-5-year cycles come about: denoting the seasonal frequency by $f_{0}=1/\text{yr}$ and the intrinsic frequency by $f_{1}$ as above, the combined effect of the two time-scales can be qualitatively understood by considering the product of two sinusoidal signals with different frequency:
$\sin(f_{0}t)\sin(f_{1}t)=\tfrac{1}{2}\left(\cos((f_{0}-f_{1})t)+\cos((f_{0}+f_{1})t)\right)$
the resulting signal thus contains a time-scale which is slower than either of the two component time-scales. In our case, the frequency $f_{0}-f_{1}$ corresponds to a period of 5.1 years (95%CI: [4.8, 5.5]). It should be noted that $5\text{yr}/43.46\text{wk}=6$ to within 1%, while 5 years is of course also an integer multiple of the period of the seasonal driver (1 year). This suggests that the system displays subharmonic resonance (the second mechanism for population cycling described in (43) and a well-known phenomenon in non-linear oscillatory systems (44, 45)).
Grossman remarked (46, 42) that the SIR equilibrium (w. vital dynamics) tends to be only weakly stable, since the real part of the eigenvalue (the decay rate) is much smaller than the imaginary part (the frequency). We find the same phenomenon for the SIRS system with M. pneumoniae parameters, with the ratio $|\text{Im}\{\lambda\}/\text{Re}\{\lambda\}|\approx 14$ (with $\lambda$ being the relevant eigenvalue of the Jacobian matrix for the linearized system).
A simple Fourier transform of the 2010-2020 (pre-COVID) data yields a dominant period of 5.1 years, while a Lomb-Scargle (LS) periodogram (47), which has finer frequency resolution, places the peak at 4.9 years. For the MpCF-based 1958-1995 time series, the LS periodogram shows the annnual signal as being the strongest, which makes sense given the observation of (2) that the multiannual outbreak pattern was somewhat overtaken by annual dynamics during the late 1970’s and early 1980’s. However, the strongest multiannual peak for the whole period (1958-1995) is at 4.1yrs (LS)/4.2yrs (Fourier). If the 1958-1975 period is viewed in isolation, the strongest peak (eclipsing the annual signal as well) is at 4.6yrs (LS)/4.3yrs (Fourier). See Supplementary Fig. S2 for Fourier and LS spectra of the incidence time series.
Demographic stochasticity
Spurred by the observation that the deterministic model will always eventually decay into annual dynamics, it is natural to ask how a departure from determinism affects the dynamics. Demographic stochasticity is a simple and ubiquitous source of noise in population dynamics (including epidemiological dynamics) (48, 49). This type of stochasticity occurs in finite populations and is fundamentally due to
1) Individuals being discrete (as opposed to the real numbers $S$, $I$ and $R$ of a typical ordinary differential equation-based model), and
2) The presence of probability rates such as $\mu$, $\beta$, $\gamma$ and $\delta$, which implies stochasticity in the timing of birth/death, infection, recovery and immune waning in finite populations.
Both of these effects can be taken into account by stochastic simulation, as may be implemented using the Doob-Gillespie algorithm (details in Materials and Methods) (50). With such a seasonally-forced stochastic SIRS model with vital dynamics, we find that multi-year cycles may indeed be sustained by demographic noise, suggesting that the system exhibits a stochastic resonance (42). However, clear multiannual cycles can be sustained by demographic stochasticity only in populations of less than approximately 2 million individuals, see Fig. 2. Simple demographic stochasticity alone is thus unlikely to be the source of sustained multiannual cycles in M. pneumoniae.
Figure 2: Demographic stochasticity excites multi-year cycles in smaller populations. Doob-Gillespie simulations of a seasonally forced SIRS model parametrized with Mycoplasma pneumoniae parameters as determined by Bayesian fit to 2010-2025 data. A) In small populations (initial population size $N_{0}=20,000$), demographic stochasticity leads to noisy multi-year cycles. B) In intermediate populations ($N_{0}=250,000$), multi-year cycles are prominent and fairly regular. C) In larger populations ($N_{0}=2\times 10^{6}$), annual dynamics dominates as the impact of demographic stochasticity is smaller.
Environmental stochasticity
We go on to consider the effects of extra-demographic or environmental stochasticity. Random fluctuations in the transmission rate $\beta(t)$ will be the primary source of environmental stochasticity considered, but we will also explore the effects of more generic exogenous shocks to the system.
As shown in Fig. 3A, random fluctuations in the transmission rate tends to stimulate approximately 5-year cycles for a wide range of moderate noise levels. When stochasticity is very slight, the annual cycle dominates, and when environmental fluctuations are very strong, the epidemic dynamics lack predictable patterns and become dominated by large but rare outbreaks, appearing as a Fourier spectrum with all frequencies represented and no discernible peaks. Details of the implementation of stochastic transmission rates, which we base on (51), can be found in Materials and Methods. It should be noted that even when the annual peak is the highest, a subdominant peak with e.g. 50% of the peak power of the annual signal will lead to a prevalence timeseries that appears very clearly multiannual. The averaging process of Fig. 3A leads to broader and flatter multiannual peaks than those found in individual spectra, and for this reason Fig. 3B shows the spectra of a few individual realizations in which the peaks at approx. 5 years stand out.
While the 5-year cycle is the most typical multiannual cycle at moderate noise levels (given M. pneumoniae parameters), other cycles are excited as well, as shown in Fig. 3, with approximately three-year as well as 6-7 year cycles being somewhat common at moderate-to-high noise levels ($\sigma>0.2$).
In Fig. 4, we explore the immediate impact of a singular perturbation (a “kick”) to the system (52). The perturbation is implemented as either a discrete change in the infected population (Fig. 4A-C) or the susceptible population (Fig. 4D). Since only a single “kick” is given, the system will only transiently settle into a multiannual cycle and we thus measure the strength of the 5-year cycle during the 20 years immediately following the perturbation. Fig. 4A shows that the timing of the perturbation matters, with a kick during the low-transmission part of the season most likely to excite the multi-year cycle. In other words, a kick of a certain size (whether in the positive or negative direction) has a larger effect if it occurs when circulation is already low. However, Fig. 4B reveals that this seasonal dependence is completely accounted for by the relative magnitude of the perturbation, i.e. the change in incidence divided by the current incidence at that time. If plotted as a function of this relative magnitude, $\Delta I/I$, the strength of the excited cycle shows only a very faint seasonal dependence.
The annual as well as the (transient) five-year cycles are shown in the $(S,I)$ configuration space (with $R=1-S-I$ suppressed) in Fig. 4C. The rate at which $I(t)$ and $S(t)$ changes can be judged by the spacing between the colored markers, and the color of the markers themselves indicate the time of year. As is evident, the five-year cycle can be quite complex. The perturbation that initiated the five-year cycle is visible as the approximately vertical line connecting the annual and multiannual cycles.
Figure 3: Moderate environmental noise stimulates multiannual cycles, especially approximately 5-year cycles. A) Average Fourier spectrum across 1560 realizations (for each noise level indicated in the legend) of the prevalence $I(t)$ of the seasonal SIRS model with parameters determined by a fit to 2010-2025 data from Denmark. The noise source is random fluctuations in the transmission rate $\beta(t)$, implemented as Gamma-distributed multiplicative noise with mean $1$ and variance $\sigma^{2}$ (see Methods and (51)) The noise level shown is the value of $\sigma$. Each analyzed simulation timeseries was 100 years, and a 1000 year “transient” was discarded. B) Fourier spectra of four individual realizations at low to moderate noise levels.
Figure 4: One-off exogeneous perturbations (“kicks”) transiently excite multiannual cycles. A) After a discrete change in prevalence of magnitude $\Delta I$, the system may temporarily settle into a 5-year cycle. The 5-year cycle is particularly prominent if the perturbation happens near the time of lowest prevalence. The color indicates the power of the 5-year cycle in the Fourier spectrum, relative to the highest peak. The spectrum is based on the first 20 years post-perturbation. The red cross indicates the perturbation implemented in panel D. B) If the magnitude of the perturbation $\Delta I$ is proportional to the current prevalence $I(t)$, the prominence of the 5-year cycle ceases to depend on time of year. The red cross again indicates the perturbation implemented in panel D. C) Annual and 5-year cycles depicted in $(S,I)$ configuration space. The marker color denotes the time of year, and markers are one week apart. The approximately vertical line starting from the red cross shows the exogenous “kick” which causes a transition from an annual (thick yellow line) to a 5-year cycle (thin purple line). D) Perturbations to $S(t)$, rather than the prevalence.
Beyond the post-COVID rebound of M. pneumoniae
Although the deterministic seasonal SIRS model captures the 2010-2025 dynamics well, predicting or forecasting M. pneumoniae incidence may be challenging. We know that long-term predictions using the determinstic model are limited by its tendency to revert to annual dynamics. In the short term, however, a deterministic model may still be appropriate for forecasting. Fig. 5A shows the trajectory obtained by continuing the seasonal SIRS model for another 10 years (with parameters from the fit to 2010-2025 data). Supplementary Fig. S1 shows that the system in fact exhibits transient chaos, with a maximal local/instantaneous Lyapunov exponent (LLE) (53) that is mostly positive. A Lyapunov exponent $\lambda$ measures sensitivity to initial conditions – this also includes the tendency of the system to amplify ($\lambda>0$) or buffer ($\lambda<0$) perturbations from e.g. noise (54). The arithmetic mean value of the greatest LLE during the first two years is approx. $0.02/\text{wk}$. In practical terms, a separation between two trajectories will have grown by almost a factor of $10$ after the first two years of deterministic dynamics. LLEs were computed using the QR method described in (55).
Moving beyond deterministic dynamics, the effects of environmental stochasticity on predictions are shown in 5B. Here, we assume a moderately high level of environmental stochasticity ($\sigma=0.15$, compare with the amplitude of seasonality at $0.19$) and explore the range of predicted outcomes. The fit to empirical data is assumed known, and noise is only introduced for the prediction part of the simulations. While sensitivity to stochastic effects may complicate medium-term predictions (note the spread between realizations in 5B from 2027 onward), some short-term features are relatively robust even at this noise level. Predictions agree that 2025 is likely to be a moderate-to-high incidence year, exemplifying how deterministic models yield valuable short-term forecasts even in transiently chaotic epidemiological systems. While long-term predictions using the deterministic model are limited by its decay into annual dynamics, the presence of stochasticity limits extended prediction not only by causing variability in the incidence during any given year, but also by enabling mode-hopping, i.e. transitions between annual and multiannual cycles (56, 49), as exemplified in Supplementary Fig. S6. That the predictions for the years immediately following the large NPI-induced perturbation agree well between the stochastic and deterministic simulations fits well with the general observation that restorative forces are stronger far from equilibrium leading the stochastic model to (temporarily) behave more like the deterministic model (42).
Figure 5: Prediction is hard. A) Prediction using deterministic model, parameters fitted to 2010-2025 data. B) Range of predictions given identical assumptions, but including noise in transmission ($\sigma_{\beta}=0.15$, 100 realizations). Noise can significantly affect long- and medium-term predictions, but all realizations agree that 2025 will see significant M. pneumoniae activity.
Limitations
In fitting the 2010-2025 time series, we used the PCR positivity rate as an incidence proxy, while we used the number of positive tests for the 1958-1995 data (total number of tests were not on record). Neither of these measures are perfect incidence proxies, and they suffer in distinct ways from biases introduced by varying testing intensities. If testing intensity is increased, this leads to an increase in the number of positive tests, while it leads to a decrease in the test positivity rate (since the denominator – the total number of tests – is increased). It may be possible to enhance fitting further by accounting for fluctuations in testing intensity when such data is available, something we have not attempted to do, beyond simple variance stabilization. Furthermore, the fidelity of the 1958-1995 data series is affected by both the (somewhat imperfect) test characteristics of the cold agglutination and complement fixation tests (since the samples considered positive are those which were positive in both of these tests (2)). We have not taken age structure into account (57), nor have we considered spatial effects which might enhance the impact of demographic stochasticity (58).
Discussion and conclusion
The origins of non-trivial cycles in ecological systems and the extent to which stochasticity plays a role in shaping them are long-standing questions in ecology, including epidemiology. In (59), the authors used wavelet analysis to identify 3-4 year cycles in pertussis across 12 countries and noted that the periodicity was transient in nature, in line with results of (60, 61). While the origins of the multiannual cycles of M. pneumoniae are generally regarded as unclear (62, 21), a few different mechanisms of varying complexity have been presented in the literature. Zhang et al. (63) suggested a combination of strain dynamics, contact network structure and waning immunity as the explanation, while Omori et al. (64) explored three different possible explanations: school-term forcing, a low-dispersion distribution of immunity duration, and interference between M. pneumoniae serotypes. They find that, of these three explanations, a less variable duration of immunity (compared with the typically assumed exponential waning) best captures the generally observed patterns in that it allows for incidence oscillations.
Recently, the authors of (26) used a TSIR model to describe the post-COVID rebound of M. pneumoniae in Denmark as well as four UN regions (Europe, Asia, the Americas, and Oceania) but found that they required at least a 90% reduction in the transmission rate due to non-pharmaceutical interventions in order to fit the data.
We find that intrinsic, noisy dynamics is sufficient to explain the complex cycles, offering a parsimonious model for the dynamics of M. pneumoniae. Earn et al. (65) demonstrated the existence of a rich landscape of stable cycles in the seasonally driven SEIR model at high transmission rates (basic reproductive numbers in excess of approx. 4). In the current case of M. pneumoniae however, multiannual cycles are unstable and require stochasticity or perturbations to be sustained and appear to be a result of stochastic resonance.
With the COVID-19 NPIs phased out, another major – if less abrupt – change in M. pneumoniae epidemiology is on the horizon. Macrolide resistance has been growing in recent years, albeit with large geographical differences (66), and it is currently unclear how this will affect epidemiological dynamics. More generally, there is a need for comparative studies of the factors governing the diversity of complex cycles observed globally for pathogens such as M. pneumoniae and parvovirus B19, which may also facilitate co-estimation of environmental sources of stochasticity.
Table 1: Estimated and assumed parameters from fit to 2010-2025 Mycoplasma pneumoniae surveillance data from Denmark.
Parameter Description Estimated value [95% HPD]
$\beta_{0}$ Mean transmission rate $0.72/\text{wk}$ $[0.68;0.79]$
$a$ Amplitude of seasonality $0.186$ $[0.174;0.197]$
$\delta$ Rate of waning of immunity $1/(11.2\text{yr})$ $[1/14.4;1/9.8]$
Parameter Description Assumed value
$\mu$ Birth/death rate $1/(80\text{yr})$
$\gamma$ Recovery rate $1/(2.5\text{wk})$
[…]
The seasonal SIRS model
The seasonally-forced SIRS model with vital dynamics and constant population size is given by
$\frac{\text{d}S}{\text{d}t}$ $=\mu-\beta(t)S(t)I(t)+\delta R(t)-\mu S(t),$ (1)
$\frac{\text{d}I}{\text{d}t}$ $=\beta(t)S(t)I(t)-(\gamma+\mu)I(t),$ (2)
$\frac{\text{d}R}{\text{d}t}$ $=\gamma I(t)-(\delta+\mu)R(t),$ (3)
where the instantaneous transmission rate $\beta(t)$ is a non-negative and periodic function with a period of one year, i.e. $\beta(t+1\text{yr})=\beta(t)$. The remaining parameters are the birth/death rate $\mu$, recovery rate $\gamma$ and immunity waning rate $\delta$. Noise in the transmission rate $\beta(t)$ is implemented as described in (51); when numerically integrating the above set of differential equations, we draw multiplicative noise from a Gamma distribution with shape $1/\sigma^{2}$ and scale $\sigma^{2}$. The parameter $\sigma$ is the noise level referred to in Fig. 3.
Linearization of the autonomous system.
The analogous autonomous system to equations (1)-(3) is obtained by substituting a constant transmission rate $\beta$ in place of $\beta(t)$ (with $\beta$ equal to the mean of its time-varying counterpart). If released away from equilibrium, $I(t)$ performs an underdamped oscillation before the system settles into an equilibrium configurations $(S^{*},I^{*},R^{*})$, where:
$S^{*}=\frac{\gamma+\mu}{\beta},\leavevmode\nobreak\ \leavevmode\nobreak\ I^{*}=\frac{(\beta-\mu-\gamma)(\delta+\mu)}{(\gamma+\delta+\mu)\beta},\leavevmode\nobreak\ \leavevmode\nobreak\ R^{*}=\frac{(\beta-\gamma-\mu)\gamma}{(\gamma+\delta+\mu)\beta}.$
Linearizing the system around this equilibrium and applying a small perturbation, the decay rate and characteristic frequency are given by the eigenvalues of the Jacobian matrix:
$J=\left(\begin{array}[]{ccc}\frac{\gamma\delta-\beta(\delta+\mu)}{\gamma+\delta+\mu}&-\gamma-\mu&\delta\\ \frac{(\delta+\mu)(\beta-\gamma-\mu)}{\gamma+\delta+\mu}&0&0\\ 0&\gamma&-\delta-\mu\\ \end{array}\right)$
We leave out the analytic expressions for the eigenvalues since they are exceedingly long and can be determined from the above matrix by standard methods. With the parameters fitted to the Danish 2010-2025 time series, the eigenvalues are $\lambda_{1}=-2.4\times 10^{-4}$ and $\lambda_{\pm}=-0.0017\pm 0.023i$. Since all real parts are strictly negative, the equilibrium is asymptotically stable.
Bayesian fitting of time series
We fit the seasonal SIRS model as described above using the probabilistic programming language Stan (67, 68), with a sinusoidal seasonally forced transmisson rate $\beta(t)=\beta_{0}\xi(t)(1+a\sin(2\pi t/\text{year}+\phi))$, to positivity rates given in the 2010-2025 data set. $\xi(t)$ is the NPI multiplier which is fixed to 1 until March 2020 and otherwise estimated. The priors used were:
  • $\beta_{0}\sim N[0;5/T]$ where $N[x;y]$ indicates a Gaussian (normal) distribution with mean $x$ and standard deviation $y$, while $T=1/\gamma$ is the recovery time (2.5wk).
  • $\text{NPI multipliers }\xi(t)\sim N[0.8;0.3]$
  • $\log(\rho)\sim N[-2;0.5]$
  • $S(t=0)\sim N[0.3;0.3]$
  • $\log(x_{I})\sim N[-4;2]$ such that $I(t=0)=x_{I}S(t=0)$ (the stick-breaking method, ensuring $S(t=0)+I(t=0)\leq 1$)
  • $\phi\sim U[0;2\pi]$ where $U[x;y]$ indicates the uniform distribution with support on the interval $[x,y]$
  • $a\sim N[0;0.10]$
  • $\delta\sim N[1/(20\text{yr});1/(10\text{yr})]$
  • $\rho x_{\text{data}}\sim N[x_{\text{model}};\sigma_{\text{obs}}]$
  • $\sigma_{\text{obs}}\sim N[0;0.1]$
[…]
Doob-Gillespie simulations
The stochastic simulations of Fig. 2 are based on the following system of transition rates:
$r[\emptyset\to S]$ $=\mu N,$
$r[S\to\emptyset]$ $=\mu S,$
$r[I\to\emptyset]$ $=\mu I,$
$r[R\to\emptyset]$ $=\mu R,$
$r[S\to I]$ =$=\begin{cases}\beta(t)SI/N&\text{if }N>0,\\ 0&\text{otherwise},\end{cases}$
$r[I\to R]$ $=\gamma I,$
$r[R\to S]$ $=\delta R,$ (4)
with the time-dependent transmission rate given by $\beta(t)=\beta_{0}(1+a\sin(2\pi t/\text{year}+\phi))$. Denoting the above rates by $r_{i}$, $i\in\{1,\dots,7\}$, the Doob-Gillespie algorithm (50) thus prescribes iteratively propagating the state of the system by drawing a uniformly distributed random number $r_{1}\in[0,1]$, determining the time to next event $\tau=-(\sum_{i}r_{i})^{-1}\log(r_{1})$ and the nature of the event by drawing another random uniform number $r_{2}\in[0,1]$ and determining the smallest $j$ for which $\sum_{i=1}^{j}a_{i}>r_{2}\sum_{i=1}^{7}a_{i}$. Event (or “reaction”) $j$ is then performed and the time is incremented by $\tau$. The procedure is then reiterated until the desired final time is reached.
[…]
We thank the Statens Serum Institut (SSI) for providing Danish surveillance data. BFN acknowledges financial support from the Carlsberg Foundation (grant no. CF23-0173). SWP is a Peter and Carmen Lucia Buck Foundation Awardee of the Life Sciences Research Foundation. MHJ is supported by the Novo Nordisk Foundation, grants NNF20OC0064978 and NNF24OC0089788. BTG would like to acknowledge support from Princeton Catalysis and Princeton Precision Health.
[…]
CODE AND DATA.
All code and data used in this manuscript may be accessed at https://github.com/BjarkeFN/MycoplasmaDynamics .